# Uncertainty quantification in fine-tuned LLMs using LoRA ensembles

### v1

@@ Abstract
Fine-tuning large language models can improve task specific performance, although a general understanding of what the fine-tuned model has learned, forgotten and how to trust its predictions is still missing. We derive principled uncertainty quantification for fine-tuned LLMs with posterior approximations using computationally efficient low-rank adaptation ensembles. We analyze three common multiple-choice datasets using low-rank adaptation ensembles based on Mistral-7b, and draw quantitative and qualitative conclusions on their perceived complexity and model efficacy on the different target domains during and after fine-tuning. In particular, backed by the numerical experiments, we hypothesise about signals from entropic uncertainty measures for data domains that are inherently difficult for a given architecture to learn.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) learns conditional distributions of vocabularies useful for generative tasks, text classification and code completion. Trained on a corpus of sequences produced by humans, such as natural language text and program source code, these models have shown remarkable capabilities (Bubeck et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023). As the applications of LLMs continue to explode, informed use of these models hinge on a proper understanding of the uncertainty of their output (Huang et al., 2023; Kuhn et al., 2023; Malinin & Gales, 2021; Ren et al., 2023).
To align an LLM towards particular needs, such as answering factual questions and instructions, a common approach is to fine-tune the model using a specialised dataset targetting human interaction (Ouyang et al., 2022). Fine-tuning involves additional training of a pre-trained LLM on a smaller dataset generated with humans (Zhong et al., 2021) or by other LLMs (Peng et al., 2023; Zhang et al., 2023). By fine-tuning, the model adapts its parameters to better capture the nuances, vocabulary, and style of the target domain(Peng et al., 2023).
Even full-model fine-tuning often requires orders of magnitude less training time, but still incur the same computational complexity and memory usage. Prior studies have shown that pre-trained language models can effectively adapt to specific tasks within smaller parameter subspaces (Li et al., 2018; Aghajanyan et al., 2020), indicating an inherently low rank during the fine-tuning process. To further reduce the training time and computational complexity of fine-tuning, a common method used is Low-Rank Adaptation (LoRA) (Hu et al., 2021). As one member of the family of methods typically referred to as parameter efficient fine-tuning (PEFT), LoRA effectively reduces the number of parameters requiring training by keeping the pre-trained weights unchanged and integrating low-rank trainable matrices into each layer of the LLM transformer architecture.
@@
We contribute to addressing these questions by analyzing the uncertainty in LLMs that have been fine-tuned using LoRA for answering multiple-choice question answers (QAs). These tasks involve definite single-token target labels, which significantly simplifies the analysis while still requiring the model to have a thorough understanding of context. We explore how uncertainty estimates, divided into epistemic model uncertainty and aleatoric data uncertainty, can be utilized to observe changes in the models knowledge regarding the presented data. We use predictive entropy and mutual information to quantify uncertainty, the former containing both aleatoric and epistemic contributions whereas the latter is solely epistemic. These entropic uncertainty measures are calculated for a Bayesian posterior derived from an ensemble of LLMs fine-tuned with LoRA.
We use the pre-trained Mistral-7b (Jiang et al., 2023) model as a prior for the weight distribution, fine-tuning on the CommonsenseQA (CQA) (Talmor et al., 2019), and the Social Sciences as well as STEM components from MMLU (Hendrycks et al., 2021) multiple-choice QA datasets. By quantifying the evolution of knowledge using predictive entropy and mutual information, we show how these measures can be used to reason about complexity of the dataset, and expected model efficacy on the target domain. Code is available (Balabanov & Linander, 2024).
@@ 2 Contributions
  • We show how uncertainty quantification with entropic measures using the ensemble posteriors can be used to reason about dataset complexity and model efficacy on the target domain.
  • We analyze three common multi-choice QA datasets, CQA, MMLU STEM and MMLU Social Sciences, using LoRA ensemble posteriors derived from a pre-trained Mistral-7b model. The analyzis of the evolution of the entropic uncertainty measures during fine-tuning lets us draw quantitative conclusions on the relative complexity, out-of-distribution behaviour, and model efficacy on the different target domains.
@@ 3.1 Uncertainty quantification in LLMs
There has been significant interest in uncertainty quantification across various tasks and domains in neural networks (Gal & Ghahramani, 2015, 2016; Malinin & Gales, 2018; Ovadia et al., 2019b; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023).
This interest extends to the realm of Large Language Models (LLMs), where accurately quantifying prediction uncertainty is a key focus (Xiao et al., 2022a; Lin et al., 2022; Mielke et al., 2022; Chen & Mueller, 2023; Duan et al., 2023; Huang et al., 2023). The application of LLMs in generative tasks introduces unique challenges, notably in measuring uncertainty of the generative outputs. (Liu et al., 2019; Malinin & Gales, 2021; Kuhn et al., 2023; Lin et al., 2023). The disentanglement of uncertainty into aleatoric and epistemic was recently discussed in the context of LLMs (Hou et al., 2023). However, it was done via ensembling of the model inputs rather than the model instances, and not in the context of fine-tuning tasks.
@@ 3.2 Fine-tuning in LLMs
Fine-tuning has recently become an integral part of the LLM ecosystem where it is used to target specific tasks such as general instruction-following model, through methods like Reinforcement Learning from Human Feedback (Houlsby et al., 2019; Hu et al., 2021; Liu et al., 2019; Ding et al., 2022, 2023).
The large computational demands of training and fine-tuning LLMs have resulted in development of more efficient techniques, known as parameter-efficient fine-tuning (PEFT)(Liu et al., 2022; Ding et al., 2022, 2023; Shi & Lipani, 2024). These methods typically involve training a small number of additional parameters on top of a fixed, pre-trained LLM, with a prominent approach being the use of low-rank adapters (LoRA) for each weight matrix (Hu et al., 2021).
@@ 3.3 LLMs within Bayesian methods
Bayesian inference techniques in neural networks have a long-standing history (Denker et al., 1987; Tishby et al., 1989; Buntine & Weigend, 1991; MacKay, 1991). These methods establish a systematic approach to derive reliable and interpretable uncertainty estimates. Previous research on Bayesian language models has been primarily concentrated on the pretraining phase of language models, rather than on their fine-tuning (Tran et al., 2019; Xue et al., 2021; Cinquin et al., 2021; Zhang et al., 2018; Chen & Li, 2024).
Recent studies have begun exploring the fine-tuning of language models using a Bayesian approach. For instance, in (Fan et al., 2020) and (Zhang et al., 2021), the attention modules are sampled either from simplex-constrained attention distributions or using Bayesian Belief Networks. Neither of these studies address entropic uncertainty measures, nor do they utilize ensembling or PEFT methods for posterior approximation.
(Yang et al., 2024) employs a post-hoc Laplace approximation (Mackay, 1992) to model LoRA parameters for fine-tuning. While this study does use LoRA for fine-tuning, it does not explore entropic uncertainty measures and focuses on the posterior over LoRA parameters rather than the model as a whole, which could limit straightforward Bayesian interpretability. There are also strong indications that deep ensembles provide more accurate posteriors compared to single-model stochastic methods like Laplace and Monte Carlo dropout (Gustafsson et al., 2019; Ovadia et al., 2019a; Fort et al., 2019; Wilson & Izmailov, 2020; Dwaracherla et al., 2022; Balabanov et al., 2023).
@@ 3.4 Ensembling LLMs
Two recent studies have explored the use of ensembling in fine-tuning LLMs, specifically focusing on full model fine-tuning method where all weights are optimized (Gleave & Irving, 2022; Sun et al., 2022). This approach has large memory overhead by construction. While uncertainty quantification using variance across the ensembles is considered, their methods lack a Bayesian formalism that could provide a more grounded interpretation and understanding.
An alternative method, BatchEnsemble (Wen et al., 2020), employs a base model with modifications through multiplicative, component-specific rank-1 matrices. This method has been applied to LLMs, but in the context of pre-training rather than fine-tuning (Tran et al., 2022).
There are also recent endeavors using LoRA ensembles for fine-tuning LLMs (Wang et al., 2023; Zhai et al., 2023). Although these studies consider uncertainty quantification, they do so without employing a Bayesian framework and do not clearly separate epistemic and aleatoric components, critial for data interpretability.
@@ 4 Bayesian Deep Learning for LLMs
Given a model with parameters $\theta$, the predictive distribution $p(t^{*}|s^{*},\mathcal{D})$ can be expressed as follows
@@
The posterior parameter distribution $p(\theta|\mathcal{D})$ contains information about the epistemic uncertainty associated with model parameters $\theta$, while $p(t^{*}|s^{*},\theta)$ reflects aleatoric uncertainty inherent in the data. Thus, as any data driven model, the output of LLMs has an associated uncertainty stemming from an uncertainty about model parameters, as well as an uncertainty regarding the correct output given the input sequence.
An example of aleatoric uncertainty is given by the input sequence “The color of the sky is ”. One would expect a model trained on a large corpus of text to produce a predictive distribution with support on the colors. This posterior distribution tells us that the model is fairly certain that the next token is a color, but that there is an uncertainty about which color. The uncertainty about the color cannot be decreased by improving the training data, it is inherent in the input sequence.
Moreover, aleatoric uncertainty depends on the architecture of the model. For example, when a current generation LLM is asked a question like “The square root of 123456789 equals ”, we expect the predictive distribution to exhibit high aleatoric uncertainty. Although the answer is definitive, and the model understands the context, it might still be incapable of providing the correct answer. This limitation might be inherent in the models architecture, and cannot be resolved through further training.
On the other hand, the input sequence “Large language models are ” presents a simple conditional completion task, but we would not trust the output predictions from a model trained on a text corpus compiled before the concept was introduced. This uncertainty is not captured in the output distribution, but rather in the uncertainty of the model parameters themselves. This epistemic uncertainty can be systematically calculated within a Bayesian formalism, where it corresponds to shape of the posterior distribution of model parameters conditioned on the training data.
Separating the origin of uncertainty is important to make informed decisions about how to interpret the preditions of LLMs. Should we trust the output for a given sequence at all, and given such trust, what is the inherent uncertainty of the output?
@@ 4.1 Fine-tuning
Note that in this formulation, there is no emphasis on the origin of the pre-trained parameters. Typically, for large language models, the training procedure consists of maximum likelihood optimization combined with supervised reinforcement learning (Touvron et al., 2023b) and the specific details of the training procedure are often not publicly disclosed. In our formulation, we do not need to delve into these origins and instead formally take it as prior knowledge upon which we base our fine-tuning.
@@ 4.2 Posterior approximation methods
Variational inference (Barber & Bishop, 1998; Graves, 2011; Blundell et al., 2015) offers a computationally efficient approximation method for the Bayesian posterior. It involves exploring a family of distributions $q_{\omega}(\theta)$, parameterized by $\omega$, to approximate the posterior $p(\theta|\mathcal{D})$. A prevalent approach to identify the optimal $q_{\omega}(\theta)$ in this set involves minimizing the Kullback-Leibler (KL) divergence $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))$ (Gawlikowski et al., 2021). This divergence, which measures information loss when $q_{\omega}(\theta)$ approximates $p(\theta|\mathcal{D})$, can be expressed as:
@@ 5.1 Low-Rank Adaptation
Low-Rank Adaptation (LoRA) provides a computationally efficient alternative to full-model fine-tuning by keeping the pre-trained weight matrices static and adding trainable low-rank matrices into each transformer layer, thereby decreasing the total number of parameters requiring training (Hu et al., 2021).
@@ 5.2 LoRA Deep Ensembles
We use LoRA ensembles for approximating the posterior $p(\theta|\mathcal{D}_{\text{fine-tune}})$. Ensembles of LoRA members have been introduced recently (Wang et al., 2023), here we provide a Bayesian treatment.
Deep ensembles can be explicitly interpreted within the Bayesian variational framework by assuming an ansatz in the form of a sum of sharply peaked distributions around parameter realizations $\omega_{k}$, with $k$ indexing ensemble members (Hoffmann & Elster, 2021; Balabanov et al., 2023). In this context, the variational KL loss simplifies to a sum of conventional single-member log likelihood loss terms, augmented with L2 regularization.
We adopt the deep ensemble variational inference formulation to include LoRA reparametrization. The trainable low-rank LoRA matrices $A_{k}$ and $B_{k}$ are not treated as a subset of the variational parameters $\omega_{k}$ of the deep ensemble ansatz. Rather, they are auxiliary degrees of freedom that parameterize these variational parameters. That is, we multiply matrices $A_{k}$ and $B_{k}$ to obtain $W_{k}=(W_{\text{pretrained}}+A_{k}B_{k})\subset\omega_{k}$, and $W_{k}$ is used in the deep ensemble ansatz for approximating the posterior distribution. The LoRA parametrisation helps us to efficiently and inexpensively find good values for $W_{k}\subset\omega_{k}$.

### v2

@@ Abstract
Fine-tuning large language models can improve task specific performance, although a general understanding of what the fine-tuned model has learned, forgotten and how to trust its predictions is still missing. We derive principled uncertainty quantification for fine-tuned LLMs with posterior approximations using computationally efficient low-rank adaptation ensembles. We analyze three common multiple-choice datasets using low-rank adaptation ensembles based on Mistral-7b, and draw quantitative and qualitative conclusions on their perceived complexity and balance between retained prior knowledge and domain specific adaptation during and after fine-tuning. We identify unexpected retention of acquired knowledge during fine-tuning in the overfitting regime. Code: github.com/oleksandr-balabanov/equivariant-posteriors
@@ 1 Introduction
Large language models (LLMs) learns conditional distributions of vocabularies useful for generative tasks, text classification and code completion. Trained on a corpus of sequences such as natural language text and program source code, these models have shown remarkable capabilities (Bubeck et al., 2023; Touvron et al., 2023a; b; Jiang et al., 2023). As the applications of LLMs continue to explode, informed use of these models hinge on a proper understanding of the uncertainty of their output (Huang et al., 2023; Kuhn et al., 2023; Malinin & Gales, 2021; Ren et al., 2023).
To align an LLM towards particular needs, such as answering factual questions and instructions, a common approach is to fine-tune the model using a specialised dataset targetting human interaction (Ouyang et al., 2022). Fine-tuning involves additional training of a pre-trained LLM on a smaller dataset generated with humans (Zhong et al., 2021) or by other LLMs (Peng et al., 2023; Zhang et al., 2023a). By fine-tuning, the model adapts its parameters to better capture the nuances, vocabulary, and style of the target domain(Peng et al., 2023).
Even full-model fine-tuning often requires orders of magnitude less training time, but still incur the same computational complexity and memory usage. Prior studies have shown that pre-trained language models can effectively adapt to specific tasks within smaller parameter subspaces Li et al. (2018); Aghajanyan et al. (2020), indicating an inherently low rank during the fine-tuning process. To further reduce the training time and computational complexity of fine-tuning, a common method used is Low-Rank Adaptation (LoRA) (Hu et al., 2021). As one member of the family of methods typically referred to as parameter efficient fine-tuning (PEFT), LoRA effectively reduces the number of parameters requiring training by keeping the pre-trained weights unchanged and integrating low-rank trainable matrices into each layer of the LLM transformer architecture.
@@
We contribute to addressing these questions by analyzing the uncertainty in LLMs that have been fine-tuned using LoRA for answering multiple-choice question answers (QAs). These tasks involve definite single-token target labels, which significantly simplifies the analysis while still requiring the model to have a thorough understanding of context. We use predictive entropy and mutual information to quantify inherent- and model uncertainty. These entropic uncertainty measures are calculated for a Bayesian posterior derived from an ensemble of LLMs fine-tuned with LoRA.
Using the pre-trained Mistral-7b Jiang et al. (2023) model as a prior, we fine-tune on the CommonsenseQA (CQA) Talmor et al. (2019) and evaluate its performance on both CQA and the Social Sciences and STEM components of the MMLU Hendrycks et al. (2021) multiple-choice QA datasets. By quantifying the evolution of knowledge using predictive entropy and mutual information, we demonstrate how these measures can be used to reason about the balance between retained prior knowledge and acquired domain specific knowledge during and after fine-tuning.
@@ 2 Contributions
  • We show how uncertainty quantification with entropic measures can be used to reason about the balance between retained knowledge and domain specific adaptation.
@@
  • We analyze three common multi-choice QA datasets, CQA, MMLU STEM and MMLU Social Sciences, using LoRA ensemble posteriors derived from a pre-trained Mistral-7b model, and identify unexpected retention of acquired knowledge in the overfitting regime.
@@ 3.1 Uncertainty quantification in LLMs
There has been significant interest in uncertainty quantification across various tasks and domains in neural networks Gal & Ghahramani (2015; 2016); Malinin & Gales (2018); Ovadia et al. (2019b); Malinin et al. (2021); Lin et al. (2022); Kuhn et al. (2023); Lin et al. (2023).
This interest extends to the realm of Large Language Models (LLMs), where accurately quantifying prediction uncertainty is a key focus Xiao et al. (2022a); Lin et al. (2022); Mielke et al. (2022); Chen & Mueller (2023); Duan et al. (2023); Huang et al. (2023); Shorinwa et al. (2024). The application of LLMs in generative tasks introduces unique challenges, notably in measuring uncertainty of the generative outputs. Liu et al. (2019); Malinin & Gales (2021); Kuhn et al. (2023); Lin et al. (2023). The disentanglement of uncertainty into aleatoric and epistemic was recently discussed in the context of LLMs Hou et al. (2023); Ling et al. (2024). However, it was done via ensembling of the model inputs rather than the model instances, and not in the context of fine-tuning tasks.
@@ 3.2 Fine-tuning in LLMs
Fine-tuning has recently become an integral part of the LLM ecosystem where it is used to target specific tasks such as general instruction-following model, through methods like Reinforcement Learning from Human Feedback Houlsby et al. (2019); Hu et al. (2021); Liu et al. (2019); Ding et al. (2022; 2023).
The large computational demands of training and fine-tuning LLMs have resulted in development of more efficient techniques, known as parameter-efficient fine-tuning (PEFT)Liu et al. (2022); Ding et al. (2022; 2023); Shi & Lipani (2024). These methods typically involve training a small number of additional parameters on top of a fixed, pre-trained LLM, with a prominent approach being the use of low-rank adapters (LoRA) for each weight matrix Hu et al. (2021).
@@ 3.3 LLMs within Bayesian methods
Bayesian inference techniques in neural networks have a long-standing history Denker et al. (1987); Tishby et al. (1989); Buntine & Weigend (1991); MacKay (1991). These methods establish a systematic approach to derive reliable and interpretable uncertainty estimates. Previous research on Bayesian language models has been primarily concentrated on the pretraining phase of language models, rather than on their fine-tuning Tran et al. (2019); Xue et al. (2021); Cinquin et al. (2021); Zhang et al. (2018); Chen & Li (2024).
Recent studies have begun exploring the fine-tuning of language models using a Bayesian approach. For instance, in Fan et al. (2020) and Zhang et al. (2021), the attention modules are sampled either from simplex-constrained attention distributions or using Bayesian Belief Networks. Neither of these studies address entropic uncertainty measures, nor do they utilize ensembling or PEFT methods for posterior approximation.
Yang et al. (2024) employs a post-hoc Laplace approximation Mackay (1992) to model LoRA parameters for fine-tuning. While this study does use LoRA for fine-tuning, it does not explore entropic uncertainty measures and focuses on the posterior over LoRA parameters rather than the model as a whole, which could limit straightforward Bayesian interpretability. There are also strong indications that deep ensembles provide more accurate posteriors compared to single-model stochastic methods like Laplace and Monte Carlo dropout Gustafsson et al. (2019); Ovadia et al. (2019a); Fort et al. (2019b); Wilson & Izmailov (2020); Dwaracherla et al. (2022); Balabanov et al. (2023). For a more detailed discussion, see Sec. 5.4.
@@ 3.4 Ensembling LLMs
Two recent studies have explored the use of ensembling in fine-tuning LLMs, specifically focusing on full model fine-tuning method where all weights are optimized Gleave & Irving (2022); Sun et al. (2022). This approach has large memory overhead by construction. While uncertainty quantification using variance across the ensembles is considered, their methods lack a Bayesian formalism that could provide a more grounded interpretation and understanding.
An alternative method, BatchEnsemble Wen et al. (2020), employs a base model with modifications through multiplicative, component-specific rank-1 matrices. This method has been applied to LLMs, but in the context of pre-training rather than fine-tuning Tran et al. (2022).
There are also recent endeavors employing LoRA ensembles for fine-tuning LLMs Wang et al. (2023); Zhai et al. (2023). Although these studies address uncertainty quantification, they do so without adopting a Bayesian framework and do not clearly distinguish between epistemic and aleatoric components, which is critical for data interpretability. In another work Arteaga et al. (2024), a separate predictor is trained on the LoRA ensembles uncertainty metrics to detect hallucinations. Niu et al. (2024) introduces a mixture-of-expert framework (UQ4CT) to explicitly capture and calibrate model uncertainty during fine-tuning. All these works are highly relevant yet differ from ours, which provides a principled Bayesian treatment for LoRA ensembles during fine-tuning and examines the evolution of uncertainty metrics to draw conclusions about the balance between retained prior knowledge and newly acquired domain adaptation.
@@ 4 Bayesian Deep Learning for LLMs
Given a model with parameters $\theta$, the predictive distribution
@@
is obtained by marginalizing over the posterior distribution of the parameters. The posterior $p(\theta\mid\mathcal{D})$ represents the uncertainty associated with model parameters $\theta$, quantifying how likely each choice of $\theta$ is, given the observed data. For a fixed $\theta$, the term $p(t^{*}\mid s^{*},\theta)$ captures data uncertainty, reflecting the variability in the model’s output $t^{*}$ for the input $s^{*}$. By integrating over all possible parameter values $\theta$, the final predictive distribution combines both uncertainty in the parameters with the inherent variability in the data Gawlikowski et al. (2023).
@@ 4.1 Fine-tuning
Note that in this formulation, there is no emphasis on the origin of the pre-trained parameters. Typically, for large language models, the training procedure consists of maximum likelihood optimization combined with supervised reinforcement learning Touvron et al. (2023b) and the specific details of the training procedure are often not publicly disclosed. In our formulation, we do not need to delve into these origins and instead formally take it as prior knowledge upon which we base our fine-tuning.
@@ 4.2 Posterior approximation methods
Variational inference Barber & Bishop (1998); Graves (2011); Blundell et al. (2015) offers a computationally efficient approximation method for the Bayesian posterior. It involves exploring a family of distributions $q_{\omega}(\theta)$, parameterized by $\omega$, to approximate the posterior $p(\theta|\mathcal{D})$. A prevalent approach to identify the optimal $q_{\omega}(\theta)$ in this set involves minimizing the Kullback-Leibler (KL) divergence $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))$ Gawlikowski et al. (2023). This divergence, which measures information loss when $q_{\omega}(\theta)$ approximates $p(\theta|\mathcal{D})$, can be expressed as:
@@ 5.1 Low-Rank Adaptation
Low-Rank Adaptation (LoRA) provides a computationally efficient alternative to full-model fine-tuning by keeping the pre-trained weight matrices static and adding trainable low-rank matrices into each transformer layer, thereby decreasing the total number of parameters requiring training Hu et al. (2021).
@@ 5.2 LoRA Deep Ensembles
We use LoRA ensembles to approximate the posterior $p(\theta|\mathcal{D}_{\text{fine-tune}})$. Ensembles of LoRA members have been introduced recently Wang et al. (2023), here we provide a Bayesian treatment.
Deep ensembles can be explicitly interpreted in a Bayesian variational framework by assuming an ansatz consisting of a mixture of sharply peaked Gaussians centered at each parameter realization $\omega_{k}$, with $k$ indexing ensemble members Hoffmann & Elster (2021); Balabanov et al. (2023). With this ansatz, the variational objective reduces to the sum of each member’s log-likelihood term plus an L2 regularization.
For additional context, deep ensembles can alternatively be viewed as a particle-based variational inference scheme, where each member evolves under Wasserstein gradient descent to minimize the KL divergence D' Angelo & Fortuin (2021); Wild et al. (2023). Both interpretations naturally give rise to a repulsive contribution, stemming from the prior term in the KL divergence, that encourages diversity and improves the posterior approximation Balabanov et al. (2023); D' Angelo & Fortuin (2021); Wild et al. (2023). In the present work, we omit this repulsive term and defer its exploration to future research.
We build on the variational deep‐ensemble framework of Hoffmann et al.Hoffmann & Elster (2021) and Balabanov et al.Balabanov et al. (2023), extending it to incorporate the LoRA reparametrization. The trainable low-rank LoRA matrices $A_{k}$ and $B_{k}$ are not treated as a subset of the variational parameters $\omega_{k}$ of the deep ensemble ansatz. Rather, they are auxiliary degrees of freedom that parameterize these variational parameters. That is, we multiply matrices $A_{k}$ and $B_{k}$ to obtain $W_{k}=(W_{\text{pretrained}}+A_{k}B_{k})\subset\omega_{k}$, and $W_{k}$ is used in the deep ensemble ansatz for approximating the posterior distribution. The LoRA parametrisation helps us to efficiently and inexpensively find good values for $W_{k}\subset\omega_{k}$.